OpenStudy (anonymous):

7x-5=8x+4. Find x. Explain your steps.

OpenStudy (anonymous):

1.

OpenStudy (anonymous):

I said explain all steps

OpenStudy (anonymous):

First you subtract 4, and you get 7x-1=8x. then subtract 7x from 8x, and you get -1=x. so my bad, it's -1 not 1.

OpenStudy (anonymous):

That's it :)

jimthompson5910 (jim_thompson5910):

the idea is to isolate x, so we need to get all the x terms to one side and everything else to the other

jimthompson5910 (jim_thompson5910):

7x-5=8x+4 7x-5-8x=8x+4-8x ... Subtract 8x from both sides. -x-5=4 ... Combine like terms (notice how the 8x terms cancel on the right) I'll let you finish up

OpenStudy (anonymous):

okay i got x=-1

OpenStudy (anonymous):

Thanks guys

jimthompson5910 (jim_thompson5910):

no, that's not correct

jimthompson5910 (jim_thompson5910):

but I can see how you got that

OpenStudy (anonymous):

Stupid me. I know all the hard things but forgot how to do this easy problem. -_-

jimthompson5910 (jim_thompson5910):

you need to add 5 to both sides to undo the minus 5 -x-5=4 -x-5+5=4+5 ... Add 5 to both sides. -x + 0 = 9 ... Combine like terms. -x = 9 x = ??? and don't say that about yourself, we all make silly mistakes

OpenStudy (anonymous):

wait hold on this is what I did

OpenStudy (anonymous):